At 25 °C, what is the hydroxide ion concentration, [OH−], in an aqueous solution with a hydrogen ion concentration of [H+] = 2.10 × 10− 3 M?
1
Expert's answer
2020-04-27T01:45:30-0400

Solution:

Auto-ionization of water is defined as the process of formation of hydronium (H3O+) and hydroxide (OH−) ions in aqueous solutions. 

2H2O(l) ⇌ H3O+(aq) + OH−(aq)

The concentration of these ions can be related together through the auto-ionization equlibrium constant (Kw):

Kw = [H3O+] × [OH−] = 10-14


Since [H3O+] is equal to [H+], then:

[OH−] = Kw / [H+];

[OH−] = (10-14) / (2.10×10-3) = 4.76×10-12

[OH−] = 4.76×10-12 M.


Answer: 4.76×10-12 M is the hydroxide ion concentration, [OH−].